# Gravitational lensing and shadows in the toron solution of Einstein’s equations using ray tracing methods

### v1

@@ 4. Geodesics in stationary axisymmetric spacetimes
In this section we briefly summarize the approach of [8] to compute geodesics in general stationary axisymmetric spacetimes. We show how individual geodesics are computed, as well as the simulation of the apparent images of extended objects in the presence of a strongly gravitating body, such as a black hole or a gravitating disk. For details the reader is referred to [8].
4.1. Geodesics
Due to the existence of two Killing vectors
$\eta=\partial_{t}\;,\hskip 19.91684pt\tilde{\eta}=\partial_{\phi}\;,$
the general geodesic equations
(30) $\frac{d^{2}x^{a}}{ds^{2}}+\Gamma_{bc}^{a}\frac{dx^{b}}{ds}\frac{dx^{c}}{ds}=0,$
(31) $E:=-g_{ab}\eta^{a}p^{b}=fp^{t}+(fA)p^{\phi}$
(here $p^{a}=dx^{a}/ds$ are momenta),
(32) $L:=g_{ab}\tilde{\eta}^{a}p^{b}=-fAp^{t}+\Phi p^{\phi}$
where
$\Phi:=g_{\phi\phi}=\rho^{2}f-\frac{A^{2}}{f}\;.$
The Lagrangian
$\mathcal{L}=\frac{1}{2}g_{ab}p^{a}p^{b}$
is also a conserved quantity, and since we are interested in null geodesics, the Lagrangian is always zero. However, unlike Kerr spacetimes, the existence of a fourth constant of motion (analogous to the Carter constant) is not known for general stationary axisymmetric spacetimes. The momenta $p^{t}$ and $p^{\phi}$ can be found from the linear system (31), (32) once the values of integrals $E$ and $L$ are fixed.
Computing the Christoffel symbols for the metric in the Weyl-Lewis-Papapetrou form (1), the geodesic equations (30) in Weyl coordinates can be written as the following system of ODEs for the remaining 6 variables which we combine into the vector $\mathbf{y}=(t,\rho,p^{\rho},z,p^{z},\phi)\in\mathbb{R}^{6}$:
(33) $\begin{split}\frac{dt}{ds}&=\frac{1}{\rho^{2}}\left(E\Phi-(fA)L\right),\\ \frac{d\rho}{ds}&=p^{\rho},\\ \frac{dp^{\rho}}{ds}&=\frac{1}{2h}\left[-f_{\rho}(p^{t})^{2}-h_{\rho}(p^{\rho})^{2}+h_{\rho}(p^{z})^{2}+\Phi_{\rho}(p^{\phi})^{2}-2(fA)_{\rho}p^{t}p^{\phi}-2h_{z}p^{\rho}p^{z}\right],\\ \frac{dz}{ds}&=p^{z},\\ \frac{dp^{z}}{ds}&=\frac{1}{2h}\left[-f_{z}(p^{t})^{2}+h_{z}(p^{\rho})^{2}-h_{z}(p^{z})^{2}+\Phi_{z}(p^{\phi})^{2}-2(fA)_{z}p^{t}p^{\phi}-2h_{\rho}p^{\rho}p^{z}\right],\\ \frac{d\phi}{ds}&=\frac{1}{\rho^{2}}((fA)E+fL),\end{split}$
(34) $\left\{\begin{array}[]{cc}\frac{d\mathbf{y}}{ds}=F(\mathbf{y}),\\ \mathbf{y}(0)=\mathbf{y}_{0},\end{array}\right.$
where $F:\mathbb{R}^{6}\to\mathbb{R}^{6}$ is the function described by (33). Hence, given initial conditions for all 8 variables ($t_{0}$, $p^{t}_{0}$, $\rho_{0}$, $p^{\rho}_{0}$, $z_{0}$, $p^{z}_{0}$, $\phi_{0}$, $p^{\phi}_{0}$), we first obtain the conserved quantities $E$ and $L$ and find the momenta $p^{t}$ and $p^{\phi}$ from the linear system (31), (32). For the remaining 6 variables encoded by the vector $\mathbf{y}$ we express the geodesic equations in the form (34) and solve them numerically with the approach presented in [8]. The IVP is solved via the Runge-Kutta (RK) method of fourth order, which gives solutions $\mathbf{y}_{1},\mathbf{y}_{2},\ldots,\mathbf{y}_{n}$ at steps $s_{1},s_{2},\ldots,s_{n}$, for some $n\in\mathbb{N}$. Thus, the function $F$ must be computed for each stage of the RK method, but this can be computationally expensive as in the case of the metric functions of toron spacetimes, since they are non-elementary functions. However, they are piecewise smooth and the method outlined in [8] to compute them on a grid via polynomial interpolation can be used. Finally, the third constant of motion $\mathcal{L}=0$ will be used to control the accuracy of the numerical solution.
Using the geodesic equations, we will study two special types of light rays. First, we look at light rays traveling initially on a plane (the equatorial plane and a plane containing the symmetry axis). Second, in the case of Kerr and NUT spacetimes, we study light rays whose radial component $r$ in Boyer-Lindquist coordinates remain constant, meaning that they are traveling on the surface of a sphere, which we refer to as the photon sphere. Similar light rays are studied for toron spacetimes, where Boyer-Lindquist coordinates are no longer natural.
@@ 4.2. Light rays on the equatorial plane and the plane passing through the symmetry axis
To make sure the light rays initially stay on the equatorial plane, we assume that the initial conditions satisfy $z_{0}=p^{z}_{0}=0$. The deviation of the light rays from the equatorial plane is determined by the value of $dp^{z}/ds$: they will move in upwards (downwards) direction if this value is positive (negative). For spacetimes whose metric functions $f$ and $e^{2k}$ are even in $z$, such as Kerr and NUT, their derivatives with respect to $z$ vanish identically on the equatorial plane. Therefore, for such spacetimes
(35) $\left.\frac{dp^{z}}{ds}\right|_{s=0}=-\left.\frac{f\cdot\partial_{z}A}{h}\right|_{(\rho_{0},z_{0})}p^{t}_{0}\,p^{\phi}_{0}.$
@@
On the other hand, in NUT spacetimes with $\ell>0$ , the derivative $\partial_{z}A$ is strictly positive on the equatorial plane (as observed in section 2). Therefore, in the case of the NUT metric the light rays do not stay on the plane, unless they move in a purely radial direction. Otherwise, they will go in downwards (respectively upwards) direction if $p^{\phi}_{0}$ is positive (respectively negative). On the other hand, due to the lack of symmetry of the metric functions in toron spacetimes, the component $dp^{z}/ds$ of the geodesic equations cannot be simplified to the form (35) and we must rely on numerical computations in order to analyze the behavior of geodesics with initial conditions of this type.
@@
where the superscripts indicate the initial value of the metric functions. Hence, the geodesic equations (33) imply
(36) $\frac{d\phi}{ds}=\frac{1}{\rho^{2}}f\cdot f^{(0)}\cdot(A-A^{(0)})p^{t}_{0}\;.$
The component $p_{0}^{t}$ is positive for light rays chosen to be moving forward in time. Moreover, we know that the function $f$ is strictly positive outside the ergosphere. Therefore, equation (36) indicates that the clockwise or counterclockwise deviation of light rays of this type is controlled by the difference $(A-A^{(0)})$ for $s\geq 0$. For instance, if the light rays are initially moving in an inwards radial direction, they bend in counterclockwise (resp. clockwise) direction if $\partial_{\rho}A|_{s=0}$ is positive (resp. negative). This conclusion holds for any stationary axisymmetric spacetime.
@@ 4.3. Photon spheres in Kerr and light rings in NUT
For Kerr and NUT spacetimes there exist light-like geodesics traveling on spheres of fixed Boyer-Lindquist radii $r$ [32, 31, 27, 24], called the photon spheres. Such geodesics are called the fundamental photon orbits [7]. If the closed light-like geodesic in Kerr spacetime lies in the equatorial plane, it is called the light ring; in NUT spacetimes the NUT rings can be lifted from the equatorial plane. In Kerr spacetimes there are two light rings: the prograde one (the direction of photons coincides with the direction of rotation) of smaller radius $r_{min}$ and the retrograde one (the direction of photons is the opposite to with the direction of rotation) or larger radius $r_{max}$. The radii of photon spheres where photons not constrained to the equatorial plane are moving, lie between $r_{min}$ and $r_{max}$.
A necessary condition for the photon to stay on the photon sphere is $dr/ds=0$ for all $s\in\mathbb{R}$. Therefore, also $d^{2}r/ds^{2}=0$ for all $s\in\mathbb{R}$. Thus, the possible photon sphere radii $r_{p}$ in a given spacetime in the presence of rotation depends on the initial condition of the null geodesic and can be found from the system of equations $dr/ds=d^{2}r/ds^{2}=0$. Such values have been computed explicitly for a family of multiparametric spacetimes, see [14], which include Kerr and NUT as particular examples.
In the equatorial plane of NUT spacetime there is only one photon ring since the metric coefficient $g_{\phi t}$ vanishes there. Its radius depends on both the mass $m$ and the gravimagnetic mass $\ell$. From the conditions for the radii of the light ring given in [14], the Boyer-Lindquist radius of the equatorial light ring is given explicitly by the real solution of the equation
@@
(37) $r_{p}=m+2\sqrt{m^{2}+\ell^{2}}\cos\left(\frac{1}{3}\arctan(\ell/m)\right),$
@@
The shape of fundamental photon orbits not staying in the equatorial plane is not known analytically. It is even more non-trivial to determine the explicit shape or even existence of the fundamental photon orbits (i.e. photon orbits staying in a neighbourhood of a compact object) and corresponding photon surfaces in general stationary axisymmetric spacetimes (including the toron spacetimes); since no equivalent to the Carter constant is known, the geodesic equations cannot be decoupled. In order to determine numerically the fundamental photon orbits in NUT and toron spacetimes we use the ray tracing techniques discussed below.
@@ 4.4. Relativistic ray tracing
In this paper we choose fixed values of $d_{H}$ and $d_{V}$. Thus, the angular apertures and therefore, the field of vision of the observer will depend inversely on the value of the focal length $d_{L}$. Once the camera features are set, each pixel on the screen gives an initial condition for the IVP (34), which is then integrated backwards in time in order to determine the sources of light.
@@
The following are the chosen criteria to interrupt the time integration of the geodesics. The computations are stopped if:
@@
If the photon hits an object, we record the color of the point hit by such a photon and associate that to the corresponding pixel. This condition is added as an event function in Matlab. If the photon approaches the first ergosphere such that $f<\epsilon$, the color given to the corresponding pixel is black which would indicate that a ray cannot originate from any light source without passing through the ergosphere. The computations are stopped since the coordinate time will go to infinity; therefore, the ergosophere cannot be crossed by any light ray in Weyl coordinates. The condition $r>R_{\infty}$ indicates that the light ray has originated from a large sphere with radius $R_{\infty}$ concentric with the strongly gravitating object. In the sequel, we refer to such a sphere as the celestial sphere in analogy with the terminology used in astronomy. If a photon meets this condition, we assign a color to the corresponding pixel according to a predefined coloring of the celestial sphere. Since the spheres with radii $R_{c}$ and $R_{\infty}$ are chosen to be concentric (see Fig. 10), the effective horizontal and vertical angles of view $\delta_{H}$ and $\delta_{V}$ for a given $R_{\infty}$ in Minkowski spacetime are given by the formula
@@
Thus, given a camera with aperture angles $\delta^{c}_{H,V}$, the observed portion of the celestial sphere will be within angles $\delta_{H,V}$ (notice that the limit when $R_{\infty}$ tends to infinity is just $\delta_{H,V}=\delta^{c}_{H,V}$), see Fig. 12 in which $\delta_{H}=\delta_{V}=90^{\circ}$. However, this formula only holds in flat spacetimes; in gravitating spacetimes it will be an upper bound for the so-called primary image due to the bending of light, as we will observe later. Nevertheless, the deformation with respect to the contour shown in Fig. 12 (which would only correspond to the contour of the primary image in gravitating spacetimes) will indicate the type of optical effect produced by a particular class of spacetimes. Moreover, in order to observe additional properties, in this paper we add an artificial coloring to the celestial sphere following the colors shown by Fig. 11 in galactic coordinates, i.e., the celestial sphere is projected onto an ellipse using the Aitoff projection (see below). This is done in order to determine the influence of the various values parametrizing the classes of spacetimes considered here on the apparent images (for instance, $\ell$ for NUT spacetimes and $\alpha$ for toron spacetimes). Namely, we determine some of the optical effects of the gravitating object by comparing the deformation of the observed image with respect the image one would expect in a Minkowski spacetime. Finally, if the computations are interrupted because $s_{n}>S_{\max}$ we color the pixel in black, since this means that the photon has made several turns around the gravitating body without falling towards the center nor escaping to infinity. We choose a value for $S_{\max}$ such that at least the secondary copy of the celestial sphere is visible, but one could choose a sufficiently large value in order to show subsequent copies.
@@ 4.5. Shadows
We recall that in the Schwarzschild spacetime there exists a unique photon sphere. In NUT spacetimes there is a unique light ring in the equatorial plane. The Schwarzschild black holes cast therefore a circular shadow on the images seen by a distant observer, see [14]. However, photon spheres are not unique in Kerr spacetimes; their radii are in an interval $r_{\min}<r_{p}<r_{\max}$ for some positive $r_{\min}$, $r_{\max}$ with the value of $r_{p}$ depending on the initial conditions of the geodesics. Therefore, the shadow cast by the rotating black hole is not circular since the radius of the photon sphere for which the photon will fall into the horizon once it is crossed depends on the direction at which the photon approaches it with respect to the frame-dragging direction. Let us recall that in Kerr spacetimes, $r_{\min}$ corresponds to a photon in prograde motion and $r_{\max}$ to a photon in retrograde motion. Moreover, the size of the photon spheres directly determines the size of the shadow. For instance, it can be inferred from formula (37) that the size of the shadow of a Schwarzschild black hole increases monotonically with its mass $m$, while the shadow of a NUT black hole (given a fixed $m$) increases monotonically with the magnitude of its gravimagnetic mass $\ell$. The computations for toron spacetimes are less straightforward, but their image simulations together with the knowledge on Kerr and NUT spacetimes will indicate the dependence of the shadow size and shape on the imaginary mass of the toron.
@@ 5.1. Initially-parallel light rays
The component $p^{t}$ is obtained as the positive solution of $g_{\mu\nu}p^{\mu}p^{\nu}=0$, since we are considering photons moving forward in time. We will refer to these conditions as initial conditions of type (i) and (ii), respectively. Figs. 13 and 14 show light rays corresponding to initial conditions of type (i) with equispaced values $y_{0}\in[-10,10]$, while Fig. 15 shows light rays with conditions of type (ii) with equispaced values $z_{0}\in[-10,10]$. In both cases, we choose a fixed value $x_{0}=10$ and shoot a beam with 20 light rays.
In order to provide a physical explanation to observed effects in toron spacetimes which are not manifested in the NUT solution, we briefly discuss first the behaviour of light rays with initial conditions of type (i) in Kerr spacetimes. It is well known that a non-zero parameter $\varphi$ in (6) gives rise to a frame-dragging effect in the direction of the black hole rotation, which increases with the magnitude of $\varphi$ (let us recall that the angular momentum is given by $J=m^{2}\sin\varphi$). In the Schwarzschild case, pairs of light rays with the same initial conditions except for the $y$-component, in which $y_{0}>0$ for one light ray and $y^{\prime}_{0}=-y_{0}$ for the other (which induce initial conditions with $\phi^{\prime}_{0}=-\phi_{0}$ and $p^{\phi^{\prime}}_{0}=-p^{\phi}_{0}$ in Weyl coordinates), are symmetric with respect to the $x$-axis. However, due to the frame-dragging effect produced by a non-zero $\varphi$, this symmetry is lost. Indeed, Fig. 13 shows that photons in prograde motion are accelerated, while those in retrograde motion are slowed down.
@@
We continue with the same analysis in NUT spacetimes. As in the previous example, we are interested in determining how a non-zero NUT parameter affects the symmetry of these light rays when compared to those in a Schwarzschild spacetime. Moreover, it is possible to choose a vanishing mass, and thus one can analyze the purely gravimagnetic effects. In the following examples, we always choose positive $\ell$.From formula (35), we expect that light rays with conditions of type (i) go in downwards (respectively upwards) direction if $p^{\phi}_{0}$ is positive (respectively negative). This is indeed what is observed in the numerical integration of the corresponding geodesics shown Fig. 14. However, it is interesting to note that their $xy$-projection is still symmetric with respect to the $x$-axis.
@@
This frame-dragging effect in opposite directions is also observed on light rays with initial conditions of type (ii). The consequence of a non-vanishing metric function $A$ on light rays with initial conditions of this type is that $p^{\phi}$, which indicates the direction in which the light rays are deflected, will depend on the sign of $z_{0}$, as implied by the formula (36). If $m\neq 0$ and $\ell=0$, which is Schwarzschild, then the function $A$ vanishes identically, implying that the rays will stay on the same plane. If $\ell>0$, then the value of the derivative $\partial_{\rho}A|_{(\rho_{0},z_{0})}$ is positive if $z_{0}>0$ and negative if $z_{0}<0$. This explains the observed behavior in Fig. 15, in which the light rays deviate in clockwise (resp. counterclockwise) direction if $z_{0}>0$ (resp. $z_{0}<0$). The dragging effect in opposite directions increases in dependence of the gravimagnetic mass $\ell$ and it is particularly noticeable for the case $m=0$ with $\ell\neq 0$. Moreover, the $xy$ projection of two light rays with initial conditions of this type, one with $z_{0}>0$ and the other with $z^{\prime}_{0}=-z_{0}$, are symmetric with respect to the $x$-axis, meaning that such light rays are deflected with the same magnitude, but in opposite directions. This can be deduced from formula (36) and the parity of the NUT metric functions in $z$.

### v2

@@ 4. Geodesics in stationary axisymmetric spacetimes
In this section we discuss ray tracing in stationary axisymmetric vacuum spacetimes with the approach [8], see Appendix B for details. We describe the expected behavior of individual geodesics given some particular initial conditions and then explain the process to simulate the apparent images of extended objects in the presence of a strongly gravitating body, such as a black hole or a gravitating disk.
@@ 4.1. Light rays on the equatorial plane and the plane passing through the symmetry axis
To make sure the light rays initially stay on the equatorial plane, we assume that the initial conditions satisfy $z_{0}=p^{z}_{0}=0$. Considering $s$ to be an affine parameter, as used in Appendix B, the deviation of the light rays from the equatorial plane is determined by the value of $dp^{z}/ds$: they will move in upwards (downwards) direction if this value is positive (negative). For spacetimes whose metric functions $f$ and $e^{2k}$ are even in $z$, such as Kerr and NUT, their derivatives with respect to $z$ vanish identically on the equatorial plane. Therefore, for such spacetimes
(4.1) $\left.\frac{dp^{z}}{ds}\right|_{s=0}=-\left.\frac{f\cdot\partial_{z}A}{h}\right|_{(\rho_{0},z_{0})}p^{t}_{0}\,p^{\phi}_{0}.$
@@
On the other hand, in NUT spacetimes with $\ell>0$ , the derivative $\partial_{z}A$ is strictly positive on the equatorial plane (as observed in section 2). Therefore, in the case of the NUT metric the light rays do not stay on the plane, unless they move in a purely radial direction. Otherwise, they will go in downwards (respectively upwards) direction if $p^{\phi}_{0}$ is positive (respectively negative). On the other hand, due to the lack of symmetry of the metric functions in toron spacetimes, the component $dp^{z}/ds$ of the geodesic equations cannot be simplified to the form (4.1), and we must rely on numerical computations in order to analyze the behavior of geodesics with initial conditions of this type.
@@
where the superscripts indicate the initial value of the metric functions. Hence, the geodesic equations (B.4) imply
(4.2) $\frac{d\phi}{ds}=\frac{1}{\rho^{2}}f\cdot f^{(0)}\cdot(A-A^{(0)})p^{t}_{0}\;.$
The component $p_{0}^{t}$ is positive for light rays chosen to be moving forward in time. Moreover, we know that the function $f$ is strictly positive outside the ergosphere. Therefore, equation (4.2) indicates that the clockwise or counterclockwise deviation of light rays of this type is controlled by the difference $(A-A^{(0)})$ for $s\geq 0$. For instance, if the light rays are initially moving in an inwards radial direction, they bend in counterclockwise (resp. clockwise) direction if $\partial_{\rho}A|_{s=0}$ is positive (resp. negative). This conclusion holds for any stationary axisymmetric spacetime.
@@ 4.2. Photon spheres in Kerr and light rings in NUT
For Kerr and NUT spacetimes there exist light-like geodesics traveling on spheres of fixed Boyer-Lindquist radii $r$ [33, 31, 27, 24], called the photon spheres. Such geodesics are called the fundamental photon orbits in general stationary axisymmetric spacetimes if their trajectory is a closed path [7]. If this trajectory is planar, then it is referred to as a light ring. In Kerr spacetime light rings lie in the equatorial plane, while in NUT spacetimes the light rings can be lifted from the equatorial plane. In Kerr spacetimes there are two light rings: the prograde one (the direction of photons coincides with the direction of rotation) of smaller radius $r_{\min}$ and the retrograde one (the direction of photons is the opposite to with the direction of rotation) of larger radius $r_{\max}$. The radii of photon spheres where photons not constrained to the equatorial plane are moving, lie between $r_{\min}$ and $r_{\max}$.
A necessary condition for the photon to stay on the photon sphere is $dr/ds=0$ for all $s\in\mathbb{R}$. Therefore, $d^{2}r/ds^{2}=0$ for all $s\in\mathbb{R}$ is also necessary. Thus, the possible photon sphere radii $r_{p}$ in a given spacetime in the presence of rotation depends on the initial condition of the null geodesic and can be found from the system of equations $dr/ds=d^{2}r/ds^{2}=0$. Such values have been computed explicitly for a family of multiparametric spacetimes, see [14], which include Kerr and NUT as particular examples. In Schwarzschild and NUT spacetimes there exists a unique radius satisfying the above condition. Its radius depends on both the mass $m$ and the gravimagnetic mass $\ell$. From the conditions for the radii of the light ring given in [14], the Boyer-Lindquist radius of the equatorial light ring is given explicitly by the real solution of the equation
@@
(4.3) $r_{p}=m+2\sqrt{m^{2}+\ell^{2}}\cos\left(\frac{1}{3}\arctan(\ell/m)\right),$
@@
The explicit description of fundamental photon orbits not staying in the equatorial plane is not known analytically (it is known that they are light rings in NUT, but the explicit equation of the plane containing it is not). It is even more non-trivial to determine the explicit shape or even existence of the fundamental photon orbits (i.e. photon orbits staying in a neighborhood of a compact object) and corresponding photon surfaces in general stationary axisymmetric spacetimes (including the toron spacetimes); since no equivalent to the Carter constant is known, the geodesic equations cannot be decoupled and thus, the Boyer-Lindquist coordinates are no longer natural. In order to determine numerically the fundamental photon orbits in NUT and an approximation thereof in toron spacetimes we use the ray tracing techniques discussed below.
@@ 4.3. Relativistic ray tracing
In this paper we choose fixed values of $d_{H}$ and $d_{V}$. Thus, the angular apertures and therefore, the field of vision of the observer will depend inversely on the value of the focal length $d_{L}$. Once the camera features are set, each pixel on the screen gives an initial condition for the initial value problem –IVP– (B.5), which is then integrated backwards in time in order to determine the sources of light.
@@
The following are the chosen criteria to terminate the time integration of the geodesics, namely, the computations are stopped if:
  • 1.
@@
1. If the photon hits an object, we record the color of the point hit by such a photon and associate that to the corresponding pixel. This condition is added as an event function in Matlab.
2. If the photon approaches the first ergosphere such that $f<\epsilon$, the color given to the corresponding pixel is black, which would indicate that a ray reaching that pixel cannot originate from any light source without passing through the ergosphere. The computations are stopped since the coordinate time will go to infinity; therefore, the ergosphere cannot be crossed by any light ray in Weyl coordinates.
3. The condition $r>R_{\infty}$ indicates that the light ray has originated from a large sphere with radius $R_{\infty}$ concentric with the strongly gravitating object. In the sequel, we refer to such a sphere as the celestial sphere in analogy with the terminology used in astronomy. If a photon meets this condition, we assign a color to the corresponding pixel according to a predefined coloring of the celestial sphere.
4. Finally, if the computations are interrupted because $s_{n}>S_{\max}$ we color the pixel in black, since this means that the photon has made several turns around the gravitating body without falling towards the center nor escaping to infinity. We choose a value for $S_{\max}$ such that at least the secondary copy of the celestial sphere is visible, but one could choose a sufficiently large value in order to show subsequent copies.
4.3.1. Angle of view and coloring of the celestial sphere
Since the spheres with radii $R_{c}$ and $R_{\infty}$ are chosen to be concentric (see Fig. 8), the effective horizontal and vertical angles of view $\delta_{H}$ and $\delta_{V}$ for a given $R_{\infty}$ in Minkowski spacetime are given by the formula
@@ 4.3.1. Angle of view and coloring of the celestial sphere
Thus, given a camera with aperture angles $\delta^{c}_{H,V}$, the observed portion of the celestial sphere will be within angles $\delta_{H,V}$ (notice that the limit when $R_{\infty}$ tends to infinity is just $\delta_{H,V}=\delta^{c}_{H,V}$), see Fig. 10 in which $\delta_{H}=\delta_{V}=90^{\circ}$. However, this formula only holds in flat spacetimes; in gravitating spacetimes it will be an upper bound for the so-called primary image due to the bending of light, as we will observe later. Nevertheless, the deformation with respect to the contour shown in Fig. 10 (which would only correspond to the contour of the primary image in gravitating spacetimes) will indicate the type of optical effect produced by a particular class of spacetimes. Moreover, in order to observe additional properties, in this paper we add an artificial coloring to the celestial sphere following the colors shown by Fig. 9 in galactic coordinates, i.e., the celestial sphere is projected onto an ellipse using the Aitoff projection (see below). This is done in order to determine the influence of the various values parametrizing the classes of spacetimes considered here on the apparent images (for instance, $\ell$ for NUT spacetimes and $\alpha$ for toron spacetimes). Namely, we determine some of the optical effects of the gravitating object by comparing the deformation of the observed image with respect the image one would expect in a Minkowski spacetime.
@@ 4.4. Shadows
We recall that in the Schwarzschild and NUT spacetimes there exists a unique photon sphere. These black holes cast therefore a circular shadow on the images seen by a distant observer, see [14]. However, photon spheres are not unique in Kerr spacetimes; their radii are in an interval $r_{\min}<r_{p}<r_{\max}$ for some positive $r_{\min}$, $r_{\max}$ with the value of $r_{p}$ depending on the initial conditions of the geodesics. Therefore, the shadow cast by the rotating black hole is not circular since the radius of the photon sphere for which the photon will fall into the horizon once it is crossed depends on the direction at which the photon approaches it with respect to the frame-dragging direction. Let us recall that in Kerr spacetimes, $r_{\min}$ corresponds to a photon in prograde motion and $r_{\max}$ to a photon in retrograde motion. Moreover, the size of the photon spheres directly determines the size of the shadow. For instance, it can be inferred from formula (4.3) that the size of the shadow of a Schwarzschild black hole increases monotonically with its mass $m$, while the shadow of a NUT black hole (given a fixed $m$) increases monotonically with the magnitude of its gravimagnetic mass $\ell$. The computations for toron spacetimes are less straightforward, but their image simulations together with the knowledge on Kerr and NUT spacetimes will indicate the dependence of the shadow size and shape on the imaginary mass of the toron.
@@ 5.1. Initially-parallel light rays
The component $p^{t}$ is obtained as the positive solution of $g_{\mu\nu}p^{\mu}p^{\nu}=0$, since we are considering photons moving forward in time. We will refer to these conditions as initial conditions of type (i) and (ii), respectively. Figs. 11 and 12 show light rays corresponding to initial conditions of type (i) with equispaced values $y_{0}\in[-10,10]$, while Fig. 13 shows light rays with conditions of type (ii) with equispaced values $z_{0}\in[-10,10]$. In both cases, we choose a fixed value $x_{0}=10$ and shoot a beam with 20 light rays.
In order to provide a physical explanation to observed effects in toron spacetimes which are not manifested in the NUT solution, we briefly discuss first the behavior of light rays with initial conditions of type (i) in Kerr spacetimes. It is well known that a non-zero parameter $\varphi$ in (2.6) gives rise to a frame-dragging effect in the direction of the black hole rotation, which increases with the magnitude of $\varphi$ (let us recall that the angular momentum is given by $J=m^{2}\sin\varphi$). In the Schwarzschild case, pairs of light rays with the same initial conditions except for the $y$-component, in which $y_{0}>0$ for one light ray and $y^{\prime}_{0}=-y_{0}$ for the other (which induce initial conditions with $\phi^{\prime}_{0}=-\phi_{0}$ and $p^{\phi^{\prime}}_{0}=-p^{\phi}_{0}$ in Weyl coordinates), are symmetric with respect to the $x$-axis. However, due to the frame-dragging effect produced by a non-zero $\varphi$, this symmetry is lost. Indeed, Fig. 11 shows that photons in prograde motion are accelerated, while those in retrograde motion are slowed down.
@@
We continue with the same analysis in NUT spacetimes. As in the previous example, we are interested in determining how a non-zero NUT parameter affects the symmetry of these light rays when compared to those in a Schwarzschild spacetime. Moreover, it is possible to choose a vanishing mass, and thus one can analyze the purely gravimagnetic effects. In the following examples, we always choose positive $\ell$.From formula (4.1), we expect that light rays with conditions of type (i) go in downwards (respectively upwards) direction if $p^{\phi}_{0}$ is positive (respectively negative). This is indeed what is observed in the numerical integration of the corresponding geodesics shown Fig. 12. However, it is interesting to note that their $xy$-projection is still symmetric with respect to the $x$-axis.
@@
This frame-dragging effect in opposite directions is also observed on light rays with initial conditions of type (ii). The consequence of a non-vanishing metric function $A$ on light rays with initial conditions of this type is that $p^{\phi}$, which indicates the direction in which the light rays are deflected, will depend on the sign of $z_{0}$, as implied by the formula (4.2). If $m\neq 0$ and $\ell=0$, which is Schwarzschild, then the function $A$ vanishes identically, implying that the rays will stay on the same plane. If $\ell>0$, then the value of the derivative $\partial_{\rho}A|_{(\rho_{0},z_{0})}$ is positive if $z_{0}>0$ and negative if $z_{0}<0$. This explains the observed behavior in Fig. 13, in which the light rays deviate in clockwise (resp. counterclockwise) direction if $z_{0}>0$ (resp. $z_{0}<0$). The dragging effect in opposite directions increases in dependence of the gravimagnetic mass $\ell$ and it is particularly noticeable for the case $m=0$ with $\ell\neq 0$. Moreover, the $xy$ projection of two light rays with initial conditions of this type, one with $z_{0}>0$ and the other with $z^{\prime}_{0}=-z_{0}$, are symmetric with respect to the $x$-axis, meaning that such light rays are deflected with the same magnitude, but in opposite directions. This can be deduced from formula (4.2) and the parity of the NUT metric functions in $z$.